Homer is considering a project with cash inflows of $950 in year 1, $900 in year 2, $1,000 in year 3 respectively. The project has a required discount rate of 11 percent and an initial cost of $2,100. What is the discounted payback period?
A. 0-1 years
B. 1-2 years
C. 2-3 years
D. more than 3 years
Answer #1

Present value = 950 / (1 + 0.11)1 = 855.8559

Present value = 900 / (1 + 0.11)2 = 730.46019

Present value = 1000 / (1 + 0.11)3 = 731.191381

Cumulative cash flow for year 0 = -2100

Cumulative cash flow for year 1 = -2100 + 855.8559 = -1,244.1441

Cumulative cash flow for year 2 = -1,244.1441 + 730.46019 = -513.68391

Cumulative cash flow for year 3 = -513.68391 + 731.191381 = 217.51

513.68391 / 731.191381 = 0.70

Discounted payback period = 2 + 0.7 = 2.7 years

C. 2-3 years